How are the classical laws of conservation of energy and conservation of mass modified by modern relativity?

1
Expert's answer
2021-04-19T07:25:25-0400

In classical mechanics, the law of energy conservation was a fundamental law. The mass was a constant that doesn't change. In relativity, the mass is not constant anymore, with the connection "E=mc^2" mass can be converted into other kinds of energy.

"E^2 =(pc)^2 +(mc^2)^2"